OpenStudy (anonymous):

Quilt squares are cut on the diagonal to form triangular quilt pieces. The hypotenuse of the resulting triangles is 26 inches long. What is the side length of each piece? (1 point) 13 13 sqrt 2 13 sqrt 3 26 sqrt 2

OpenStudy (austinl):

Pythagorean Theorem is this, \(c^2=a^2+b^2\) But because it is a square the side lengths are the same meaning that a=b \(c^2=2a^2\) You know c=26 Do you think you could solve from there?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

13